A fair die is tossed until the first 1 appears. Let x equal the number of tosses required, x = 1,2,3,... You are to receive (.5)x dollars if the first appears on the x-th toss. What is the expected amount that you will receive? Hint: infinite geometric series.
Answer #1

P(1 on dice) = 1/6

E(x) = 1 / P(1 appearing)

= 1 / (1/6)

= 6

E(ax) = a*E(x)

amount recieved = 0.5*x

E(amount)

= E(0.5x)

= 0.5*E(x)

= 0.5*6

= $3

(please UPVOTE)